I know equations but I have no idea how to solve word problems of equations. Please help me!

The admission fee at a small fair is $1.50 for children and $4.00 for adults. On a certain day, 2200 people enter the fair and $5050 is collected. How many children and how many adults attended?

Answer by Earlsdon(6294)

Let C = the number of children at $1.50 each.
The number of adults is 2200 - C at $4.00 each.
The total collected from the sum of these two groups is $ 5050.00
Now you can set up the equation:
(C)($1.50) + (2200-C)($4.00) = $5050.00 Simplify and solve for C.
1.5C + 8800 - 4C = 5050
-2.5C + 8800 = 5050 Subtract 8800 from both sides.
-2.5C = -3750 Divide both sides by -2.5
C = 1500
The number of children attended was 1500
The number of adults attended was 2200 - 1500 = 700